One positive number is four times another number. If half the smaller number is added to the larger number, the result is 18. What are the numbers?
Also how did you solve it, please show equations

Answer by mananth(16949)

let one number be x
the other number = 4x
.
x/2 + 4x=18
LCM =2
x+8x / 2 =18
9x=36
x=4
The other number will be 4x = 4*4 =16